Calculate the area expansivity and volume expansivity of a brass of area 120m and 120m respectively that assumes a new area and volume of 130m and 130m respectively when the linear expansivity is 0.00005 through a temperature of 50 degrees Celsius
1
Expert's answer
2020-06-21T20:09:18-0400

Area expansivity = 2 * Linear expansivity

 "\u00df = 2 * \u221d"

Area expansivity = 2 * 0.00005 = 0.0001


Volume expansivity = 3 * Linear expansivity

 "\u03d2 = 3*\u221d"

Volume expansivity = 3 * 0.00005 = 0.00015